artlover03:

please help!! the function f and g are defied by *will post* which equation is equivalent to **will post**

artlover03:

1 attachment
artlover03:

@jhonyy9 @snowflake0531

snowflake0531:

Do you get what you are supposed to do?

snowflake0531:

Like, I know you posted this because you don't know, but like, do you know, like, generally, what the equation is supposed to be, and how it would be simplified

artlover03:

no i dont, functions arent really my thing

snowflake0531:

Let's take it apart f(2x) multiplied by g(-2x) divided by 2 f(x) = x^2 g(x) = 2x

snowflake0531:

So, let's first do f(2x) Literally, this just means, in f(x) = x^2, put x, as 2x So, what is it?

snowflake0531:

Substitute x as 2x

artlover03:

... f(2 x 2x)?

snowflake0531:

._. f(x) = x^2 Now, it has become f(2x) = ....? In, x^2 , substitute x as 2x

artlover03:

2x^2??

snowflake0531:

Yep, so it's 2x^2 times g(-2x) over 2 Now, we do g(-2x) We know that g(x) = 2x So, in ^^^, plug in x as -2x

artlover03:

2x^2?

snowflake0531:

g(x) = 2x g(-2x) =....

snowflake0531:

x is now -2x Put that into 2x

artlover03:

-4x^2? i suck at this

snowflake0531:

First of all, I said wrong earlier, the previous one is 4x^2, not 2x^2

snowflake0531:

Secondly relook x=-2x g(x) = 2x Now that x is -2x You have to substitute it 2(-2x)

artlover03:

-4x^2..

snowflake0531:

2 times -2x

artlover03:

-4x

snowflake0531:

Yes Now multiply f(2x) with g(-2x) 4x^2 times -4x

artlover03:

-16x^3

snowflake0531:

Now what is -16x^3 divided by 2

artlover03:

-8x^3

snowflake0531:

There's your answer~ Although I have a question lol Do you understand how I got to that, or were you only answering what I asked you?

artlover03:

no i got it; took my a while to catch on :)

snowflake0531:

Oh, okay, that's nice then

artlover03:

thank you :))

snowflake0531:

Yw~!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 2x^2?? \(\color{#0cbb34}{\text{End of Quote}}\) no - not is right if f(x) = x^2 so f(2x) = (2x)^2 what not equal 2x^2 @snowflake0531 more ATTENTION please !!!

snowflake0531:

I said it was wrong afterwards

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 Yep, so it's 2x^2 times g(-2x) over 2 Now, we do g(-2x) We know that g(x) = 2x So, in ^^^, plug in x as -2x \(\color{#0cbb34}{\text{End of Quote}}\) not is right

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 First of all, I said wrong earlier, the previous one is 4x^2, not 2x^2 \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 Yep, so it's 2x^2 times g(-2x) over 2 Now, we do g(-2x) We know that g(x) = 2x So, in ^^^, plug in x as -2x \(\color{#0cbb34}{\text{End of Quote}}\) here you said yep - what mean yes

snowflake0531:

But afterwards, I said that I was wrong\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 \(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 First of all, I said wrong earlier, the previous one is 4x^2, not 2x^2 \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 First of all, I said wrong earlier, the previous one is 4x^2, not 2x^2 \(\color{#0cbb34}{\text{End of Quote}}\) ohh here

snowflake0531:

Yea, the above was an oops